OpenStudy (anonymous):

Find the coordinates of the other endpoint when you are given the midpoint and one of the endpoints. Answer in terms of a decimal rounded to the nearest tenth or an improper fraction in simplest form. P1 M P2 (3, 5) (-2, 0) ( , ) helpppp

OpenStudy (anonymous):

please teach me step by step

OpenStudy (anonymous):

midpoint of x = \(\Large \frac{x_1 + x_2}{2}\)

OpenStudy (anonymous):

midpoint of y = \(\Large \frac{y_1 + y_2}{2}\)

OpenStudy (anonymous):

okay...midpoint of x means x coordinate of the midpoint...what is the x coordinate of your midpoint?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

now...what is the x coordinate of one of the endpoints?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so... \[\large -2 = \frac{3 + x_2}{2}\] got that? i just substituted

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

:D yes!

OpenStudy (anonymous):

so i assume you can do the rest?

OpenStudy (anonymous):

then i do the same for y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!!!!

OpenStudy (anonymous):

you're welcome